# Local Geometric and Transport Properties of Networks that are Generated from Hyperuniform Point Patterns

### v1

@@ Abstract
Hyperuniformity, which is a type of long-range order that is characterized by the suppression of long-range density fluctuations in comparison to the fluctuations in standard disordered systems, has emerged as a powerful concept to aid in the understanding of diverse natural and engineered phenomena. In the present paper, we harness hyperuniform point patterns to generate a class of disordered, spatially embedded networks that are distinct from both perfectly ordered lattices and uniformly random geometric graphs. We refer to these networks as hyperuniform-point-pattern-induced (HuPPI) networks, and we compare them to their counterpart Poisson-point-pattern-induced (PoPPI) networks. By computing the local geometric and transport properties of HuPPI networks, we demonstrate how hyperuniformity imparts advantages in both transport efficiency and robustness. Specifically, we show that HuPPI networks have systematically smaller total effective resistances, slightly faster random-walk mixing times, and fewer extreme-curvature edges than PoPPI networks. Counterintuitively, we also find that HuPPI networks simultaneously have more negative mean Ollivier–Ricci curvatures and smaller global resistances than PoPPI networks, indicating that edges with moderately negative curvatures need not create severe bottlenecks to transport. Moreover, HuPPI networks are consistently more robust under both random edge removals and curvature-based targeted edge removals, maintaining larger connected components for larger fractions of removed edges than their PoPPI counterparts. We also demonstrate that the network-generation method strongly influences these properties and in particular that it often overshadows differences that arise from underlying point patterns. These results collectively demonstrate potential advantages of hyperuniformity in network design and motivate further theoretical and experimental exploration of HuPPI networks.
@@ I Introduction
From transportation infrastructure [4] and power grids [5] to neuronal networks [6] and protein-interaction networks [7, 8], spatial constraints play a fundamental role in shaping network structure and function. The geometric arrangement of nodes and the associated distances between them often constrain connectivity (e.g., by influencing measures like spatial strength centrality [9]), transport, and robustness [10]. Accordingly, there is considerable interest in understanding how different spatial point patterns, which range from uniformly random arrangements to highly correlated patterns, affect network properties after one uses such patterns to generate a network [2]. Indeed, the analysis of geometric networks that arise from point patterns is an important topic in the field of stochastic geometry [11]. One way to analyze the properties of such networks is through geometric data analysis and the examination of discrete curvature on networks [12]. Notions of network curvature have useful physical interpretations. For example, Devriendt et al. [13, 14] introduced a resistance-based notion of graph curvature that links effective resistance, commute times, and mixing behavior in networks. In the present paper, we work with classical Ollivier–Ricci curvature (ORC) in spatially embedded graphs (i.e., networks), and we highlight the broader theme that curvature statistics are closely related to transport efficiency. We use ORC instead of Forman–Ricci curvature [15] (which is another popular type of network curvature) because the definition of ORC is based on optimal transport, which aligns with our focus on network transport properties like total effective resistance and random-walk mixing times.
@@
Researchers have begun to explore how hyperuniformity manifests in spatially embedded networks [19, 20, 21]. One approach is to use hyperuniform or nonhyperuniform point patterns to determine the locations of the nodes of a graph and to then assign edges between nodes using some geometric construction (e.g., a Delaunay or Voronoi tessellation). Maher et al. [19] demonstrated that certain tessellation-based networks partially inherit hyperuniformity properties of the underlying point pattern. Newby et al. [20, 21] illustrated that Voronoi and Delaunay tessellations that are generated from hyperuniform point patterns have narrower distributions of cell areas and reduced local density correlations than networks that one constructs from Poisson-distributed points.
@@
In the present paper, we investigate the properties that HuPPI networks inherit from the underlying hyperuniformity of their point patterns. Much is known about the transport properties of hyperuniform two-phase media [22, 17, 23], yet there is little understanding of the transport properties of HuPPI networks. Our aim is observational: we build HuPPI and Poisson-point-pattern-induced (PoPPI) networks and study whether and how the properties of the ensuing network families differ from each other.
@@
Three lessons emerge. First, total effective resistances (TER) [24] and random‑walk mixing times [25], which are global measures of transport in networks, are smaller in HuPPI networks than in PoPPI networks. This is consistent with findings on two-phase hyperuniform materials [18], where the specific microstructure geometry and topology are the primary determinants of effective transport properties. The method of network generation establishes a structure’s fundamental topology (e.g., triangle-rich versus tree-like structures), which impacts baseline transport efficiency more strongly than the variations that arise due to underlying point patterns. However, the method of network generation (e.g., Delaunay tessellations, Voronoi tessellations, Delaunay-centroidal tessellations, or Gabriel graphs) [2, 18] has a much larger impact than the underlying point pattern on these global measures. Second, we observe that the differences that are consistent across network-generation methods are not driven directly by suppressed density fluctuations. Instead, they stem from the way that hyperuniformity biases local edge geometry, thereby narrowing the ORC distribution and reducing the frequency of extreme bottleneck edges. We compute ORC [26, 27], which is a discrete analogue of the Ricci curvature from geometry [28], to determine locally how “bottleneck-like” an edge is by comparing the optimal-transport cost that is required to transform the probability distribution of the neighborhood of one of its attached nodes into the probability distribution of the neighborhood of its other attached node. When the neighborhoods of two connected nodes share many common neighbors, the optimal-transport cost is low (i.e., the associated curvature is positive), whereas a lone bridge-like edge has a high transport cost (i.e., an associated negative curvature) [26, 27]. Third, the TERs and random-walk mixing times of HuPPI networks are smaller in magnitude than those of PoPPI networks, despite often having more negative mean curvatures. This counterintuitive result arises because hyperuniformity suppresses extreme curvature outliers. PoPPI networks possess a few edges with large-magnitude negative curvatures that act as bottlenecks and inflate the TER. By contrast, HuPPI networks do not have many edges with large-magnitude negative curvatures and instead have many edges with small-magnitude negative curvatures, which do not create severe flow constraints.
@@ II Hyperuniformity
A point pattern is hyperuniform [17] if the number variance satisfies
@@ III.2 Network generation
There are numerous models of random tessellations [31]. In our study, we construct four types of planar networks (and associated tessellations) from the 2D point patterns that we described in Sec. III.1. For each point pattern $P=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{N}\}\subset\mathbb{R}^{2}$, we generate four families of networks.
@@ IV.1 Total effective resistance (TER) $\mathcal{R}_{\mathrm{tot}}$
In Figs. 1 and 2, we illustrate how normalized TER $\mathcal{R}_{\mathrm{norm}}$ varies with the disorder strength $a$ and the network size $n$. Except for the Delaunay networks, we observe that the HuPPI networks have smaller TERs than PoPPI networks, reflecting the former’s more uniform coverage in space that suppresses large voids between points. By contrast, Poisson point patterns can generate extended spatial regions with no nodes or edges. In an undirected network, when there are such regions, current (or some other quantity) that flows between many origin–destination node pairs has to detour along longer paths, and the resulting larger pairwise effective resistances accumulate to inflate the TER $\mathcal{R}_{\mathrm{tot}}$.
The Voronoi and Delaunay-centroidal networks tend to have larger $\mathcal{R}_{\mathrm{norm}}$ values than the Gabriel and Delaunay networks because the tree-like connections in the former two types of networks limit their numbers of redundant paths. In Delaunay, Delaunay-centroidal, and Voronoi networks, the total number of edges is determined solely by the system size, so the edge count is the same for all network realizations. By contrast, different Gabriel graphs with the same system size have different numbers of edges. Consequently, Gabriel graphs have noticeably higher variabilities in $\mathcal{R}_{\mathrm{norm}}$ (see Figs. 1 and 2) than the other three types of networks.
@@ VI Robustness of networks to edge removals
We examine network robustness by progressively removing edges from our networks [40]. Our results are means of 20 independent realizations of each network type. We generate 20 independent realizations of each point pattern and then construct the four types of networks from each of these realizations. Cetinay et al. [48] demonstrated that targeting edges (or nodes) based on their curvatures or resistance-based centralities is able to identify structural bottlenecks in power-transmission networks, and they illustrated that removing such edges significantly disrupts network connectivity. Taking inspiration from their results, we consider two edge-removal strategies: (1) random removal, in which we remove edges uniformly at random; and (2) targeted removal, in which we remove edges in the order of their ORC values, starting with the edges with the most negative values (i.e., the most bottleneck-like). In our targeted-removal strategy, progressively removing the most bottleneck-like edges allows us to assess how quickly a network’s connectivity degrades as one compromises the most important pathways through it.

### v2

@@ Abstract
Hyperuniformity, which is a type of long-range order that is characterized by the suppression of long-range density fluctuations in comparison to the fluctuations in standard disordered systems, has emerged as a powerful concept to aid in the understanding of diverse natural and engineered phenomena. In the present paper, we harness hyperuniform point patterns to generate a class of disordered, spatially embedded networks that are distinct from both perfectly ordered lattices and uniformly random geometric graphs. We refer to these networks as hyperuniform-point-pattern-induced (HuPPI) networks, and we compare them to their counterpart Poisson-point-pattern-induced (PoPPI) networks. By computing the local geometric and transport properties of HuPPI networks, we demonstrate how hyperuniformity imparts advantages in both transport efficiency and robustness. Specifically, we show that HuPPI networks have systematically smaller total effective resistances, slightly faster random-walk mixing times, and fewer extreme-curvature edges than PoPPI networks. Counterintuitively, we also find that HuPPI networks simultaneously have more negative mean Ollivier–Ricci curvatures and smaller total effective resistances than PoPPI networks, indicating that edges with moderately negative curvatures need not create severe bottlenecks to transport. Moreover, HuPPI networks are consistently more robust under both random edge removals and curvature-based targeted edge removals, maintaining larger connected components for larger fractions of removed edges than their PoPPI counterparts. We also demonstrate that the network-generation method strongly influences these properties and in particular that it often overshadows differences that arise from underlying point patterns. These results collectively demonstrate potential advantages of hyperuniformity in network design and motivate further theoretical and experimental exploration of HuPPI networks.
@@ I Introduction
From transportation infrastructure [4] and power grids [5] to neuronal networks [6] and protein-interaction networks [7, 8], spatial constraints play a fundamental role in shaping network structure and function. The geometric arrangement of nodes and the associated distances between them often constrain connectivity (e.g., by influencing measures like spatial strength centrality [9]), transport, and robustness [10]. Accordingly, there is considerable interest in understanding how different spatial point patterns, which range from uniformly random arrangements to highly correlated patterns, affect network properties after one uses such patterns to generate a network [2]. Indeed, the analysis of geometric networks that arise from point patterns is an important topic in the field of stochastic geometry [11]. One way to analyze the properties of such networks is through geometric data analysis and the examination of discrete curvature on networks [12]. Notions of network curvature have useful physical interpretations. For example, Devriendt et al. [13, 14] introduced a resistance-based notion of graph curvature that links effective resistance, commute times, and mixing behavior in networks. Additionally, Robertson et al. [15] established a mathematical link between effective resistance and the $1$-Wasserstein distance in Ollivier–Ricci curvature (ORC). They demonstrated that both of these quantities solve the same transport problem, but under different cost constraints. Motivated by this connection, we study classical ORC in spatially embedded graphs (i.e., networks) and highlight the broader theme that curvature statistics are closely related to transport efficiency. We use ORC instead of Forman–Ricci curvature [16] (which is another popular type of network curvature) because the definition of ORC is based on optimal transport, which aligns with our focus on network transport properties like total effective resistance and random-walk mixing times.
@@
Researchers have begun to explore how hyperuniformity manifests in spatially embedded networks [20, 21, 23]. One approach is to use hyperuniform or nonhyperuniform point patterns to determine the locations of the nodes of a graph and to then assign edges between nodes using some geometric construction (e.g., a Delaunay or Voronoi tessellation). Maher et al. [20] demonstrated that certain tessellation-based networks partially inherit hyperuniformity properties of the underlying point pattern. Other researchers [22, 21, 23] have illustrated that Voronoi and Delaunay tessellations that are generated from hyperuniform point patterns have narrower distributions of cell areas and reduced local density correlations than networks that one constructs from Poisson-distributed points.
@@
In the present paper, we investigate the properties of HuPPI networks. Such networks do not simply inherit the hyperuniformity of their underlying point patterns, but hyperuniformity does have a major impact on their properties. Much is known about the transport properties of hyperuniform two-phase media [24, 18, 25], yet there is little understanding of the transport properties of HuPPI networks. Our aim is observational: we build HuPPI and Poisson-point-pattern-induced (PoPPI) networks and study whether and how the properties of the ensuing network families differ from each other.
@@
Three lessons emerge. First, total effective resistances (TER) [26] and random‑walk mixing times [27], which are global measures of transport in networks, are smaller in HuPPI networks than in PoPPI networks. This is consistent with findings on two-phase hyperuniform materials [19], where the specific microstructure geometry and topology are the primary determinants of effective transport properties. The method of network generation establishes a structure’s fundamental topology (e.g., triangle-rich versus tree-like structures), which impacts baseline transport efficiency more strongly than the variations that arise due to underlying point patterns. However, the method of network generation (e.g., Delaunay tessellations, Voronoi tessellations, Delaunay-centroidal tessellations, or Gabriel graphs) [2, 19] has a much larger impact than the underlying point pattern on these global measures. Second, we observe that the differences that are consistent across network-generation methods are not driven directly by suppressed large length-scale density fluctuations. Instead, they stem from the way that hyperuniformity can bias local edge geometry, thereby narrowing the ORC distribution and reducing the frequency of extreme bottleneck edges. We compute ORC [28, 29], which is a discrete analogue of the Ricci curvature from geometry [30], to determine locally how “bottleneck-like” an edge is by comparing the optimal-transport cost that is required to transform the probability distribution of the neighborhood of one of its attached nodes into the probability distribution of the neighborhood of its other attached node. When the neighborhoods of two connected nodes share many common neighbors, the optimal-transport cost is low (i.e., the associated curvature is positive), whereas a lone bridge-like edge has a high transport cost (i.e., an associated negative curvature) [28, 29]. Third, the TERs and random-walk mixing times of HuPPI networks are smaller in magnitude than those of PoPPI networks, despite often having more negative mean curvatures. This counterintuitive result arises because hyperuniformity suppresses extreme curvature outliers. PoPPI networks possess a few edges with large-magnitude negative curvatures that act as bottlenecks and inflate the TER. By contrast, HuPPI networks do not have many edges with large-magnitude negative curvatures and instead have many edges with small-magnitude negative curvatures, which do not create severe flow constraints.
@@ II Hyperuniformity
A point pattern is hyperuniform [18] (which was originally termed “superhomogeneity” [31]) if the number variance satisfies
@@ III Methods
Figure 1: Visualizations of the types of networks that we construct from hyperuniform point patterns. In the top row, we show examples of Gabriel, Delaunay, Delaunay-centroidal, and Voronoi networks that we generate from a single point pattern with disorder strength $a=1$. In the bottom row, we show Delaunay-centroidal networks that we generate using disorder strengths $a\in\{0.1,0.5,1.0,2.0\}$, illustrating the transition from a near-lattice structure to a disordered structure. We generate and analyze networks with periodic boundary conditions, but we truncate edges at the boundaries in this figure for visualization purposes.
@@ III.2 Network generation
There are numerous models of random tessellations [34]. In our study, we construct four types of planar networks (and associated tessellations) from the 2D point patterns that we described in Sec. III.1. We show examples of these networks in Fig. 1. For each point pattern $P=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{N}\}\subset\mathbb{R}^{2}$, we generate four families of networks.
@@ IV.1 Total effective resistance (TER) $\mathcal{R}_{\mathrm{tot}}$
In Figs. 2 and 3, we illustrate how normalized TER $\mathcal{R}_{\mathrm{norm}}$ varies with the disorder strength $a$ and the network size $n$. Except for the Delaunay networks, we observe that the HuPPI networks have smaller TERs than PoPPI networks, reflecting the former’s more uniform coverage in space that suppresses large voids between points. By contrast, Poisson point patterns can generate extended spatial regions with no nodes or edges. In an undirected network, when there are such regions, current (or some other quantity) that flows between many origin–destination node pairs has to detour along longer paths, and the resulting larger pairwise effective resistances accumulate to inflate the TER $\mathcal{R}_{\mathrm{tot}}$.
The Voronoi and Delaunay-centroidal networks tend to have larger $\mathcal{R}_{\mathrm{norm}}$ values than the Gabriel and Delaunay networks because the tree-like connections in the former two types of networks limit their numbers of redundant paths. In Delaunay, Delaunay-centroidal, and Voronoi networks, the total number of edges is determined solely by the system size, so the edge count is the same for all network realizations. By contrast, different Gabriel graphs with the same system size have different numbers of edges. Consequently, Gabriel graphs have noticeably higher variabilities in $\mathcal{R}_{\mathrm{norm}}$ (see Figs. 2 and 3) than the other three types of networks.
@@ VI Robustness of networks to edge removals
Having established that PoPPI networks possess outlier edges with very large $\Delta\mathcal{R}_{\mathrm{tot}}(e)$ values that correspond to critical bridges (whose removal severely disrupts global transport), we now study the robustness of HuPPI and PoPPI networks. We examine network robustness by progressively removing edges from our networks [43]. Our results are means of 20 independent realizations of each network type. We generate 20 independent realizations of each point pattern and then construct the four types of networks from each of these realizations. Cetinay et al. [51] demonstrated that targeting edges (or nodes) based on their curvatures or resistance-based centralities is able to identify structural bottlenecks in power-transmission networks, and they illustrated that removing such edges significantly disrupts network connectivity. Taking inspiration from their results, we consider two edge-removal strategies: (1) random removal, in which we remove edges uniformly at random; and (2) targeted removal, in which we remove edges in the order of their ORC values, starting with the edges with the most negative values (i.e., the most bottleneck-like). In our targeted-removal strategy, progressively removing the most bottleneck-like edges allows us to assess how quickly a network’s connectivity degrades as one compromises the most important pathways through it.
@@ VII Conclusions and discussion
Our first observation is further supported by the fact that the TERs, mixing times, and mean ORCs of the HuPPI networks converge to those of the PoPPI networks as we increase disorder strength. Importantly, this convergence occurs even though the underlying point patterns remain hyperuniform at these disorder strengths. This suggests that the network structure and point-pattern hyperuniformity decouple from each other. Because our network-generation methods depend only on the local geometry of an underlying point pattern, the resulting network structure is determined by the point pattern only at short length scales. Accordingly, these networks do not inherit hyperuniformity at large lengths scales.